# Machine Learning for Structural Biology

### v1

@@ Abstract
Machine Learning-assisted directed evolution (MLDE) is a powerful tool for efficiently navigating antibody fitness landscapes. Many structure-aware MLDE pipelines rely on a single conformation or a single committee across all conformations, limiting their ability to separate conformational uncertainty from epistemic uncertainty. Here, we introduce a rank‑conditioned committee (RCC) framework that leverages ranked conformations to assign a deep neural network committee per rank. This design enables a principled separation between epistemic uncertainty and conformational uncertainty. We validate our approach on SARS-CoV-2 antibody docking, demonstrating significant improvements over baseline strategies. Our results offer a scalable route for therapeutic antibody discovery while directly addressing the challenge of modeling conformational uncertainty.
@@ 1 Introduction
Antibody therapeutics have recently become the dominant type of drug, overtaking small molecules. As a result, it is of great interest to develop methods to design antibodies to bind a given target antigen. It is the goal of the present work to produce a computational pipeline for antibody design optimization that makes use of state-of-the art existing techniques, but overcomes remaining challenges related to conformational uncertainty. It is also a goal of this work to have the algorithms be iterative in nature, so that they can in the future be naturally used in a lab-in-the-loop format, to engage directly with experimental exploration of designs.
@@
Machine learning-assisted directed evolution (MLDE) has significantly enhanced therapeutic antibody design by accelerating the exploration of vast sequence spaces and prioritizing high-affinity variants [29, 37]. In parallel, modern folding algorithms and ensemble techniques now rapidly generate accurate protein conformations for use in in silico antibody optimization [1, 22]. Typical in silico antibody binding pipelines perform an initial stage of rigid‑body docking followed by a semi‑flexible pose refinement at the side chain and backbone interface. However, large loop rearrangements are still hard to realize unless present in the starting ensemble [19]. Flexible docking alternatives can remodel paratopes, but their performance remains sensitive to input ensembles and computationally demanding. Thus, rich conformational ensembles are critical for computationally efficient antibody optimization [7]. This is particularly important in the case of SARS-CoV-2, where state switching in the receptor binding domain (RBD) [36] and variability in the heavy-chain complementarity-determining region (CDR-H3) [6, 5] complicate binding affinity prediction [1]. Although ensemble techniques may broaden structural coverage, the critical difficulty in combining these approaches with MLDE lies in disentagling uncertainty arising from binding pose selection from uncertainty in the accuracy of the model [24]. Addressing this separation is essential for antibody design, especially against novel pathogens that stretch our current understanding and may drive future outbreaks and pandemics.
@@
In protein design, structure-aware MLDE often couples surrogate models with Bayesian active learning to propose new sequences using acquisition functions such as expected improvement or upper confidence bounds that depend on both the surrogate mean and predictive uncertainty [37]. A key challenge is that predictive variance can originate from two conceptually distinct sources: epistemic uncertainty due to model limitations and aleatoric uncertainty arising from conformational heterogeneity of antibody–antigen complexes [24]. Standard ensemble approaches such as deep ensembles generally conflate these effects. Bayesian Committee Machines provide principled ways to combine estimators [33], and recent work has explored ensemble learning over conformational docking ensembles [27]. Building on these ideas, our contribution is to introduce a rank-conditioned committee that stratifies ensembles by conformation rank, so that within-rank dispersion captures epistemic variance while between-rank dispersion captures aleatoric variance, thereby enabling acquisition functions that more reliably balance exploration and exploitation in MLDE.
@@
Figure 1 illustrates the practical challenges of modeling antibodies under conformational uncertainty. In the left panel, the residue-wise confidence scores for AlphaFold2 [23] and ImmuneBuilder [1] are reported, with CDR-H3 showing significantly lower confidence relative to the framework regions. The middle panel shows the conformations predicted by AlphaFold2 and Immunebuilder. Each of these sequences are aligned and their CDR-H3 regions are colored and circled. There is a nearly exact match outside of the CDR-H3 region, but significant variance among the CDR-H3 loops themselves.
The right panel of Figure 1 connects these modeling challenges to experimental outcomes for the best performing antibodies after five rounds of biologically informed directed evolution. The experimental method is described in detail in [21]. Briefly, DNA genes for variable heavy and light-chain sequences prodcued by the AI antibody pipeline were designed in SnapGene [11], synthesized as gBlocks (IDT), and amplified and purified for direct use in cell-free protein synthesis (CFPS). CFPS was carried out with the NEB PURExpress kit. Binding of expressed single domain antibody (sdFab) constructs to the target antigen (SARS-CoV-2 Receptor Binding Domain (RBD)) was evaluated using a solution-phase flourescent reporter assay for binding ( AlphaLISA assay [8]) : CFPS products were incubated with the RBD antigen, followed by AlphaLISA donor/acceptor bead addition, and reporter luminescence was measured on a plate reader. Signals were background-subtracted, averaged across replicates, and normalized to a literature parent antibody (ID150/SC2-31). Positive controls yielded strong Alpha signals while negative controls remained at baseline, confirming assay specificity. The figure shows the results for three different antibody designs, as normalized to the parent antibody results (each design is tested in independent triplicates (shown as three colored bars) with 6 technical replicates each (shown as error bars)), and these show a trend that the improved designs do indeed tend to produce better measured binding (with the caveats of experimental variability, and that the AlphaLISA assay is not a direct measurement of equilibrium binding affinity).
@@ 2 Machine Learning Directed Evolution (MLDE)
where $S$ is the set of all developable antibody sequences, and $h(s)$ is the performance metric, which in this case is the HADDOCK3 [19] docking score between a given antibody sequence $s$ and the fixed SARS-CoV-2 antigen. The CDR-H3 region of a given sequence $s$ includes at least 30 mutable residues, each of which can take one of 20 amino acids, yielding a sequence space of size $|S|\geq 20^{30}$. Exhaustive search over such a space is infeasible. Moreover, evaluating the objective function $h(s)$ is computationally expensive, limiting the number of sequences and poses that can be directly assessed.
It is especially useful tsolve this optimization problem through an iterative evolution of the design, based on (in silico or experimental) evaluations of intermediate candidate Binding, because such methods can readily take advantage of the demonstrated power of directed evolution methodologies. In particular, MLDE addresses these challenges by replacing exhaustive search with iterative rounds of surrogate modeling and sequence selection. The success of MLDE depends critically on strategies that (i) reduce the effective search space by incorporating biologically informed constraints, and (ii) improve sampling efficiency through Bayesian active learning. This section reviews these complementary approaches and their role in enabling tractable antibody optimization.
@@ 2.2 Bayesian Active Learning
To obtain a fixed-and-low-dimensional embedding, we apply Antibody Mutagenesis-Augmented Processing (AbMAP), a framework introduced by Singh et al. [31]. AbMAP begins with representations from foundational protein language models (PLMs), which mostly encode information across conserved regions of protein sequences. To extract embedded information from antibody hypervariable CDR regions, AbMAP also generates embeddings for randomized mutants of the input sequence for comparison. By subtracting the average mutant embedding from the original, the framework isolates CDR-specific contributions, a curated embedding that is passed through a downstream transformer-based neural network to yield fixed-dimensional representations tailored to the structural and functional landscape of antibody hypervariable regions.
@@ 3.1 Ensemble Based Acquisition Maximization
The process is iterative, where each rank–conditioned committees can be retrained and new candidate libraries can be generated.
@@ 3.2 Methods
We use the SARS‑CoV‑2 antibody database [20], and build a fully autonomous, end‑to‑end in silico pipeline by integrating open‑source tools (Riot‑NA [12], ImmuneBuilder [1], PBDFixer [13], and HADDOCK3 [19]). ImmuneBuilder[1], which is specialized for antibodies, returns four predicted conformations per sequence in PDB format. Per‑residue variability across these predicted conformations is summarized via RMSD to capture local structural uncertainty. Docking is performed using HADDOCK3 [19], a data-driven platform for modeling biomolecular complexes. HADDOCK3 integrates experimental data and physicochemical parameters to generate structural models of protein-to-protein interactions. For each antibody-antigen pair, the software produces multiple docked complexes in PDB format, along with detailed scoring information that includes van der Waals, electrostatics, desolvation energy, and an overall HADDOCK binding score. The final HADDOCK binding score is a weighted sum of energy terms, designed to approximate the quality and stability of the predicted protein–protein complex. A more negative score indicates stronger binding for a given complex, which is why we employ a minimization function. HADDOCK3 groups resulting models into clusters based on structural similarity, and only top-scoring clusters are retained for downstream analysis. This procedure enables us to prioritize variants with strong predicted binding ability and plausible interaction interfaces.
@@ 4 Experimental Results
In our experiments, we compared three main approaches: (i) bioinformatics-based directed evolution, (ii) machine learning-assisted directed evolution (MLDE) with an XGBoost ensemble, and (iii) MLDE with a deep neural network (DNN) ensemble. The primary objective was to evaluate whether our methods can consistently improve the fitness landscape by identifying higher-scoring mutants. To this end, we tracked the mean score of the population across multiple runs. An upward shift in the mean score indicates that the approach is effective in discovering beneficial variants. Moreover, the magnitude of this shift reflects the efficiency of the method: the larger and faster the improvement, the more effective the strategy is at navigating the search space. Additionally, variance plays a crucial role in evaluating the algorithm, as it reflects the diversity and explorability of the search. Algorithms that achieve a large mean shift while maintaining moderate variance are prioritized, as they demonstrate both effectiveness and the ability to explore diverse regions of the solution space.
Refer to caption Refer to caption Refer to caption
Figure 3: Comparison of results. The plots illustrate three approaches evaluated in this study: Bioinformatics-based directed evolution, MLDE with an XGBoost ensemble, and MLDE with a DNN ensemble (left to right). Each histogram shows the evaluation of the initial dataset (grey) alongside 200 generated mutants, divided into two groups of 100 according to their appearance in the pipeline. The plots report the mean and variance of each round. The strongest performance is observed in the rightmost (blue) plot with the DNN ensemble, where both the mean and variance of the mutant population are highest, reflecting effective exploration and exploitation.
@@
To ensure a fair comparison, we generated 200 sequences for each method and compared them. The results are summarized in Figure 3. The mutants were partitioned into three equal, chronological batches, starting with the initial dataset, followed by two subsequent batches, each containing 100 mutants. We ensured that all hyperparameters were consistent across methods. The specific hyperparameters can be found in Appendix D.1.
@@
We start with the initial dataset, which has a mean of 69.83 and variance of 14.86. For Bioinformatics-directed evolution, the first batch has a mean of 74.87 (+7.2%) and the second batch a mean of 75.21 (+7.7%). For MLDE with XGBoost , the first batch shows a mean of 86.89 (+24.4%) and the second batch a mean of 90.47 (+29.5%). For MLDE with DNN , the first batch has a mean of 87.69 (+25.5%) and the second batch 93.95 (+34.5%).
The results show that the mean increases in all scenarios, confirming that all three algorithms are effective. However, the DNN approach not only demonstrates the highest mean shift but also maintains a reasonable variance, indicating it achieves both high exploration and exploitation.
@@ Appendix A Background
As traditional DE still incurs high costs from wet-lab evaluation, computational tools based on machine learning have emerged to replace portions of the DE workflow with in silico counterparts. Linking such tools together makes it possible to bring the protein design problem in silico and use the aforementioned ZO algorithms for protein optimization. A vast amount of software has been developed in recent years to model protein interaction in silico, circumventing the traditionally tedious processes of NMR and X-ray crystallography for determining protein complex structure [30]. This includes HADDOCK3, which incorporates ambiguous interaction restraints to compute different possible docking configurations [18, 19], and AutoDock, which specifies flexibility and rigidity in different components of the interaction [15]. PRODIGY [35] computes binding affinity based on structural contact features. We note that different software compute different measures of “good" binding—for example, while HADDOCK models physical “fit" and computes an aggregation of energy terms as its final score (estimating $\Delta H$), AutoDock and PRODIGY estimate the binding free energy ($\Delta G$).
@@
To analyze the relationship between related proteins in terms of function and evolutionary history, biologists have used sequence alignment-based approaches to identify corresponding regions of amino acids. Such approaches rely on amino acid substitution matrices, which reflect the frequencies at which amino acids undergo substitution mutations naturally in proteins over time. Standard substitution matrices include PAM [10], which captures substitution rates among similar sequences, and BLOSUM [16, 17], which detects distant relationships by considering substitution rates across dissimilar clusters of sequences. Such biologically derived substitution matrices guide lab-induced mutagenesis with real-world observations of which changes are realistic and stable, as opposed to random sampling with equal likelihood [2].
@@
As previously mentioned, it is impossible to test all possible sequences in $S$, due to the combinatorial vastness of the search space. In general, DE can be considered as an active learning technique, since new inputs (untested variants) are proposed based on known data (performance of previously tested sequences). Active learning and cluster learning methods have successfully been integrated with wet lab directed evolution to significantly improve properties like product yield and fitness [29, 37].
Moreover, zeroth-order optimization strategies like Bayesian optimization have been applied to constrained regions of antibodies to improve binding [25]. Significant progress has been made as well to design appropriate embeddings for CDR regions [31] to treat antibody sequences as vectors, which allow sequences to be better analyzed mathematically and can be incorporated into kernels for active learning techniques.
@@
When the goal is not to design a protein de novo, but rather to improve the function of a known sequence, we can take advantage of helpful starting points for optimizing $h$. Naturally, protein functions are refined throughout thousands or even millions of years of evolution, through random mutation. Scientists can compress this process in the lab with DE. In this iterative process, the scientist induces a mutation in an amino acid sequence and tests the resulting protein’s performance on a targeted function. The sequence is either discarded, if the protein’s performance is suboptimal, or introduced to a new round of mutation and evaluation [3]. Simulating Darwininan evolution in the lab allows the researcher to isolate protein function, free of additional biological constraints imposed by the organism from which the protein originated [4]. With scientific software for simulated protein design and evaluation becoming more available, in silico directed evolution promises to be a much more cost and time effective process for optimizing our function $h$ that we are trying to minimize . A description of the software used in this project and their capabilities, inputs, and outputs can be found in Data Generation Section.
@@
Directed evolution can conveniently be viewed mathematically as a discrete ZO algorithm. ZO is concerned with optimizing black-box functions without a well-behaved or easily computable gradient [26], as in our case (more details are provided in Chapter LABEL:ch:Pipeline). ZO distinguishes itself from gradient-free and derivative-free optimization methods in that it does not model the gradient $\nabla{h}$ or attempt to extend the discrete-valued function $h$ to a continuous version. Rather, it relies solely on individual evaluations of $h$.
ZO algorithms in computer science include evolutionary algorithms [34] and Bayesian optimization [14]. Evolutionary algorithms are analogous to the directed evolution procedure in a wet lab. They select new points in the domain to evaluate that are “mutated" versions of points previously found to have low objective function values. Bayesian optimization utilizes surrogate models of the objective function and an acquisition function to determine high-value points to sample next.

### v2

# Conformational Rank Conditioned Committees for Machine Learning-Assisted Directed Evolution

@@ Abstract
Machine Learning-assisted directed evolution (MLDE) is a powerful tool for efficiently navigating antibody fitness landscapes. Many structure-aware MLDE pipelines rely on a single conformation or a single committee across all conformations, limiting their ability to separate conformational uncertainty from epistemic uncertainty. Here, we introduce a rank‑conditioned committee (RCC) framework that leverages ranked conformations to assign a deep neural network committee per rank. This design enables a principled separation between epistemic uncertainty and conformational uncertainty. We validate our RCC-MLDE approach on SARS-CoV-2 antibody docking, demonstrating significant improvements over baseline strategies. Our results offer a scalable route for therapeutic antibody discovery while directly addressing the challenge of modeling conformational uncertainty.
@@ 1 Introduction
Antibody therapeutics have recently become the dominant type of drug, overtaking small molecules. As a result, it is of great interest to develop methods to design antibodies to bind a given target antigen. Here we develop a computational pipeline for antibody design optimization that makes use of state-of-the art existing techniques while explicitly addressing conformational uncertainty. The algorithms are iterative by design, so that they can in the future be naturally used in a lab-in-the-loop format, to engage directly with experimental exploration of designs.
@@
Machine learning-assisted directed evolution (MLDE) has significantly enhanced therapeutic antibody design by accelerating the exploration of vast sequence spaces and prioritizing high-affinity variants qiu2021cluster; yang2025active. In parallel, modern folding algorithms and ensemble techniques now rapidly generate accurate protein conformations for use in in silico antibody optimization abanades2023immunebuilder; jingAlphaFoldMeetsFlow2024. Typical in silico antibody binding pipelines perform an initial stage of rigid‑body docking followed by a semi‑flexible pose refinement at the side chain and backbone interface. However, large loop rearrangements are still hard to realize unless present in the starting ensemble honorato2024haddock2. Flexible docking alternatives can remodel paratopes, but their performance remains sensitive to input ensembles and computationally demanding. Thus, rich conformational ensembles are critical for computationally efficient antibody optimization bhattTeachingOldDocks2024. This is particularly important in the case of SARS-CoV-2, where state switching in the receptor binding domain (RBD) yajimaStructuralBasisReceptorbinding2024 and variability in the heavy-chain complementarity-determining region (CDR-H3) barnesStructuresHumanAntibodies2020; asarnowStructuralInsightSARSCoV22021 complicate binding affinity prediction abanades2023immunebuilder. Although ensemble techniques may broaden structural coverage, the critical difficulty in combining these approaches with MLDE lies in disentangling uncertainty arising from binding pose selection from uncertainty in the accuracy of the model kendallWhatUncertaintiesWe2017. Addressing this separation is essential for antibody design, especially against novel pathogens that stretch our current understanding and may drive future outbreaks and pandemics.
@@
In protein design, structure-aware MLDE often couples surrogate models with Bayesian active learning to propose new sequences using acquisition functions such as expected improvement or upper confidence bounds that depend on both the surrogate mean and predictive uncertainty yang2025active. A key challenge is that predictive variance can originate from two conceptually distinct sources: epistemic uncertainty due to model limitations and aleatoric uncertainty arising from conformational heterogeneity of antibody–antigen complexes kendallWhatUncertaintiesWe2017. Standard ensemble approaches such as deep ensembles generally conflate these effects. Bayesian Committee Machines provide principled ways to combine estimators trespBayesianCommitteeMachine2000, and recent work has explored ensemble learning over conformational docking ensembles mohammadiEnsembleLearningEnsemble2022. Building on these ideas, our contribution is to introduce a rank-conditioned committee that stratifies ensembles by conformation rank, so that within-rank dispersion captures epistemic variance while between-rank dispersion captures aleatoric variance, thereby enabling acquisition functions that more reliably balance exploration and exploitation in MLDE.
@@
Figure 1 illustrates the practical challenges of modeling antibodies under conformational uncertainty. In the left panel, the residue-wise confidence scores for AlphaFold2 AlphaFold2021 and ImmuneBuilder abanades2023immunebuilder are reported, with CDR-H3 showing significantly lower confidence relative to the framework regions. The middle panel shows the conformations predicted by AlphaFold2 and Immunebuilder. Each of these sequences are aligned and their CDR-H3 regions are colored and circled. There is a nearly exact match outside of the CDR-H3 region, but significant variance among the CDR-H3 loops themselves.
The right panel of Figure 1 connects these modeling challenges to experimental outcomes for the best performing antibodies after five rounds of biologically informed directed evolution. The experimental method is described in detail in huntRapidCellfreeExpression2023. Briefly, DNA genes for variable heavy and light-chain sequences prodcued by the AI antibody pipeline were designed in SnapGene snapgene, synthesized as gBlocks (IDT), and amplified and purified for direct use in cell-free protein synthesis (CFPS). CFPS was carried out with the NEB PURExpress kit. Binding of expressed single domain antibody (sdFab) constructs to the target antigen (SARS-CoV-2 Receptor Binding Domain (RBD)) was evaluated using a solution-phase flourescent reporter assay for binding ( AlphaLISA assay bielefeld2009alphalisa) : CFPS products were incubated with the RBD antigen, followed by AlphaLISA donor/acceptor bead addition, and reporter luminescence was measured on a plate reader. Signals were background-subtracted, averaged across replicates, and normalized to a literature parent antibody (ID150/SC2-31). Positive controls yielded strong Alpha signals while negative controls remained at baseline, confirming assay specificity. The figure shows the results for three different antibody designs, as normalized to the parent antibody results (each design is tested in independent triplicates (shown as three colored bars) with 6 technical replicates each (shown as error bars)), and these show a trend that the improved designs do indeed tend to produce better measured binding (with the caveats of experimental variability, and that the AlphaLISA assay is not a direct measurement of equilibrium binding affinity).
@@ 2 Machine Learning Directed Evolution (MLDE)
where $S$ is the set of all developable antibody sequences, and $h(s)$ is the performance metric, which in this case is the HADDOCK3 honorato2024haddock2 docking score between a given antibody sequence $s$ and the fixed SARS-CoV-2 antigen. The CDR-H3 region of a given sequence $s$ includes at least 30 mutable residues, each of which can take one of 20 amino acids, yielding a sequence space of size $|S|\geq 20^{30}$. Exhaustive search over such a space is infeasible. Moreover, evaluating the objective function $h(s)$ is computationally expensive, limiting the number of sequences and poses that can be directly assessed.
Rather than attempting exhaustive search, it is natural to solve this optimization problem through via iterative evolution of the design, using in silico or experimental evaluations of intermediate candidates. Such strategies leverage the demonstrated power of directed evolution methodologies. In particular, MLDE addresses these challenges by replacing exhaustive search with iterative rounds of surrogate modeling and sequence selection. The success of MLDE depends critically on strategies that (i) reduce the effective search space by incorporating biologically informed constraints, and (ii) improve sampling efficiency through Bayesian active learning. This section reviews these complementary approaches and their role in enabling tractable antibody optimization.
@@ 2.2 Bayesian Active Learning
To obtain a fixed-and-low-dimensional embedding, we apply Antibody Mutagenesis-Augmented Processing (AbMAP), a framework introduced by Singh et al. singh2025learning. AbMAP begins with representations from foundational protein language models (PLMs), which mostly encode information across conserved regions of protein sequences. To extract embedded information from antibody hypervariable CDR regions, AbMAP also generates embeddings for randomized mutants of the input sequence for comparison. By subtracting the average mutant embedding from the original, the framework isolates CDR-specific contributions, a curated embedding that is passed through a downstream transformer-based neural network to yield fixed-dimensional representations tailored to the structural and functional landscape of antibody hypervariable regions.
@@ 3.1 Ensemble Based Acquisition Maximization
The process is iterative, where each rank-conditioned committees can be retrained and new candidate libraries can be generated.
@@ 3.2 Methods
We use the SARS‑CoV‑2 antibody database hunt2023rapid, and build a fully autonomous, end‑to‑end in silico pipeline by integrating open‑source tools (Riot‑NA riot, ImmuneBuilder abanades2023immunebuilder, PBDFixer pdbfixer, and HADDOCK3 honorato2024haddock2). ImmuneBuilderabanades2023immunebuilder, which is specialized for antibodies, returns four predicted conformations per sequence in PDB format. Per‑residue variability across these predicted conformations is summarized via RMSD to capture local structural uncertainty. Docking is performed using HADDOCK3 honorato2024haddock2, a data-driven platform for modeling biomolecular complexes. HADDOCK3 integrates experimental data and physicochemical parameters to generate structural models of protein-to-protein interactions. For each antibody-antigen pair, the software produces multiple docked complexes in PDB format, along with detailed scoring information that includes van der Waals, electrostatics, desolvation energy, and an overall HADDOCK binding score. The final HADDOCK binding score is a weighted sum of energy terms, designed to approximate the quality and stability of the predicted protein–protein complex. A more negative score indicates stronger binding for a given complex, which is why we employ a minimization function. HADDOCK3 groups resulting models into clusters based on structural similarity, and only top-scoring clusters are retained for downstream analysis. This procedure enables us to prioritize variants with strong predicted binding ability and plausible interaction interfaces.
Our RCC-MLDE Approach.
The RCC-MLDE framework runs in sequential rounds. At iteration $n$, we maintain a labeled set $D_{n}=\{(x_{i},\{y_{i,r}\}_{r=1}^{R})\}$ of all sequences $s_{i}$ that have been evaluated so far, where $x_{i}=\iota(s_{i})$ is the AbMAP embedding and $y_{i,r}$ is the HADDOCK3 docking score for each ImmuneBuilder conformational rank $r$. From the current population, we generate a candidate library $C_{n}$ via bioinformatics-guided single- and double-site substitutions and parental crossovers, then filter these sequences for basic biological feasibility (Riot-NA) and structural plausibility. For each rank $r$, we construct a rank-specific dataset $D_{n}^{(r)}=\{(x_{i},y_{i,r})\}$ and train an ensemble (“committee”) of surrogate models, either fully connected DNNs or XGBoost regressors, on $D_{n}^{(r)}$. Given a candidate embedding $x\in C_{n}$, each committee produces a rank-conditional mean and epistemic variance, which we aggregate across ranks either with a standard ensemble acquisition (baseline MLDE) or with the rank-conditioned committee acquisition defined in Section 3.1 that separately tracks epistemic and conformational uncertainty. We then score all $x\in C_{n}$, select the top $B$ sequences under the chosen acquisition, run ImmuneBuilder and HADDOCK3 to obtain new labels $\{y_{r}(x)\}_{r=1}^{R}$, and form $D_{n+1}$ by appending these measurements. There is no static train/test split; at each iteration models are trained on $D_{n}$ with bootstrapping and evaluated only on unseen candidates in $C_{n}$, and newly evaluated sequences are used for training only in subsequent rounds. We depict this pipeline in Appendix D, and our experimental results summarize the distribution of docking scores of the new antibodies.
@@ 4 Experimental Results
In our experiments, we compared three main approaches: (i) bioinformatics-based directed evolution, (ii) machine learning-assisted directed evolution (MLDE) with an XGBoost ensemble, and (iii) MLDE with a deep neural network (DNN) ensemble. The primary objective was to evaluate whether our methods can consistently improve the fitness landscape by identifying higher-scoring antibody variants. To this end, we tracked the mean score of the population across multiple runs. An upward shift in the mean score indicates that the approach is effective in discovering beneficial variants. Moreover, the magnitude of this shift reflects the efficiency of the method: the larger and faster the improvement, the more effective the strategy is at navigating the search space. Additionally, variance plays a crucial role in evaluating the algorithm, as it reflects the diversity and explorability of the search. Algorithms that achieve a large mean shift while maintaining moderate variance are prioritized, as they demonstrate both effectiveness and the ability to explore diverse regions of the solution space.
Refer to caption Refer to caption
(a) Directed Evolution (b) Baseline MLDE with DNN
Refer to caption Refer to caption
(c) RCC MLDE with XGBoost (b) RCC MLDE with DNN
Figure 2: Comparison of results for (a) bioinformatics-based DE (b) MLDE with a DNN ensemble (c) RCC-MLDE with XGBoost and (d) RCC-MLDE with a DNN ensemle. Each histogram shows the evaluation of the initial dataset (grey) alongside 200 generated antibodies, divided into two learning batches of 100 variants. The mean and variance of each batch is reported.
@@
To ensure a fair comparison, we generated 200 sequences for each method and compared them. The results are summarized in Figure 2. The antibody variants were partitioned into three equal, chronological batches, starting with the initial dataset, followed by two subsequent batches, each containing 100 mutants. We ensured that all hyperparameters were consistent across methods. The specific hyperparameters can be found in Appendix D.2.
@@
We start with the initial dataset, which has a mean of 69.83 and variance of 14.86. For Bioinformatics-directed evolution, the first batch has a mean of 74.87 (+7.2%) and the second batch a mean of 75.21 (+7.7%). For MLDE with DNN , the first batch has a mean of 87.69 (+25.5%) and the second batch 93.95 (+34.5%). For RCC MLDE with XGBoost, the first batch shows a mean of 86.89 (+24.4%) and the second batch a mean of 90.47 (+29.5%). For RCC MLDE with DNN, the first batch has a mean of 84.85 (+21.5%) and the second batch 92.02 (+31.8%).
Consider sequences in the extreme right tail of the predicted docking-score distribution in Figure 2. With a single surrogate trained on pooled conformations, some of these extremes can have high mean scores even though their conformations disagree strongly, so apparent performance depends heavily on which pose is realized. Empirically, RCC yields populations with high mean docking scores and fewer extreme right-tail points, so top-ranked candidates are strong performers without being dominated by single, idiosyncratic conformations, providing a more conservative and reproducible basis for downstream screening.
RCC introduces a modest additional cost: instead of a single committee, we train one committee per rank, so the cost scales linearly in the number of ranks $R$. In our experiments we use $R=4$ ImmuneBuilder conformations per sequence, which is the default samples returned. Computational time remained a small fraction of the overall docking time (which is dominated by HADDOCK3).
@@ 5 Conclusion and Outlook
Acknowledgments
This work was supported in part by the National Science Foundation under grant DMS-1925919.
@@ Appendix A Background
As traditional DE still incurs high costs from wet-lab evaluation, computational tools based on machine learning have emerged to replace portions of the DE workflow with in silico counterparts. Linking such tools together makes it possible to bring the protein design problem in silico and use the aforementioned ZO algorithms for protein optimization. A vast amount of software has been developed in recent years to model protein interaction in silico, circumventing the traditionally tedious processes of NMR and X-ray crystallography for determining protein complex structure [ritchie2008recent]. This includes HADDOCK3, which incorporates ambiguous interaction restraints to compute different possible docking configurations [honorato2021structural, honorato2024haddock2], and AutoDock, which specifies flexibility and rigidity in different components of the interaction [goodsell2021autodock]. PRODIGY [xue2016prodigy] computes binding affinity based on structural contact features. We note that different software compute different measures of “good" binding—for example, while HADDOCK models physical “fit" and computes an aggregation of energy terms as its final score (estimating $\Delta H$), AutoDock and PRODIGY estimate the binding free energy ($\Delta G$).
@@
To analyze the relationship between related proteins in terms of function and evolutionary history, biologists have used sequence alignment-based approaches to identify corresponding regions of amino acids. Such approaches rely on amino acid substitution matrices, which reflect the frequencies at which amino acids undergo substitution mutations naturally in proteins over time. Standard substitution matrices include PAM [PAM], which captures substitution rates among similar sequences, and BLOSUM [henikoff1992amino, henikoff1993performance], which detects distant relationships by considering substitution rates across dissimilar clusters of sequences. Such biologically derived substitution matrices guide lab-induced mutagenesis with real-world observations of which changes are realistic and stable, as opposed to random sampling with equal likelihood [adenot1999peptides].
@@
As previously mentioned, it is impossible to test all possible sequences in $S$, due to the combinatorial vastness of the search space. In general, DE can be considered as an active learning technique, since new inputs (untested variants) are proposed based on known data (performance of previously tested sequences). Active learning and cluster learning methods have successfully been integrated with wet lab directed evolution to significantly improve properties like product yield and fitness [qiu2021cluster, yang2025active].
Moreover, zeroth-order optimization strategies like Bayesian optimization have been applied to constrained regions of antibodies to improve binding [khan2022antbo]. Significant progress has been made as well to design appropriate embeddings for CDR regions [singh2025learning] to treat antibody sequences as vectors, which allow sequences to be analyzed more easily in a mathematical framework and incorporated into kernels for active learning techniques.
@@
When the goal is not to design a protein de novo, but rather to improve the function of a known sequence, we can take advantage of helpful starting points for optimizing $h$. Naturally, protein functions are refined throughout thousands or even millions of years of evolution, through random mutation. Scientists can compress this process in the lab with DE. In this iterative process, the scientist induces a mutation in an amino acid sequence and tests the resulting protein’s performance on a targeted function. The sequence is either discarded, if the protein’s performance is suboptimal, or introduced to a new round of mutation and evaluation [arnold1993engineering]. Simulating Darwininan evolution in the lab allows the researcher to isolate protein function, free of additional biological constraints imposed by the organism from which the protein originated [arnold1996directed]. With scientific software for simulated protein design and evaluation becoming more available, in silico directed evolution promises to be a much more cost and time effective process for optimizing our function $h$ that we are trying to minimize . A description of the software used in this project and their capabilities, inputs, and outputs can be found in Data Generation Section.
@@
Directed evolution can conveniently be viewed mathematically as a discrete ZO algorithm. ZO is concerned with optimizing black-box functions without a well-behaved or easily computable gradient [liu2020primer], as in our case (more details are provided in Chapter LABEL:ch:Pipeline). ZO distinguishes itself from gradient-free and derivative-free optimization methods in that it does not model the gradient $\nabla{h}$ or attempt to extend the discrete-valued function $h$ to a continuous version. Rather, it relies solely on individual evaluations of $h$.
ZO algorithms in computer science include evolutionary algorithms [wierstra2014natural] and Bayesian optimization [frazier2018tutorial]. Evolutionary algorithms are analogous to the directed evolution procedure in a wet lab. They select new points in the domain to evaluate that are “mutated" versions of points previously found to have low objective function values. Bayesian optimization utilizes surrogate models of the objective function and an acquisition function to determine high-value points to sample next.